Assume that the crust temperature increases by 30k for eack kilometer below earths's surface where the temperature is 300k. Calculate what the temperature would be at Earth's center if the rate of increase were the same below the crust as in it. The distance from the crust to the earth's center is 6371km: Assume that the crust temperature increases by 30k for eack kilometer below earths's surface where the temperature is 300k. Calculate what the temperature would be at Earth's center if the rate of increase were the same below the crust as in it. The distance from the crust to the earth's center is 6371km
Answer by ankor@dixie-net.com(4531)

Assume that the crust temperature increases by 30k for each kilometer below
earth's surface where the temperature is 300k. Calculate what the temperature
would be at Earth's center if the rate of increase were the same below the crust
as in it. The distance from the crust to the earth's center is 6371km
:
Let d = distance in km
:
T = 300 + 30d
:
T = 300 + 30(6371)
:
T = 300 + 191130
:
T = 191,430k; pretty hot!